OpenStudy (anonymous):

How would I solve? Can I have an example? The price of a notebook has dropped to 3.10 today. Yesterday's price was 3.35 . Find the percentage decrease.

jimthompson5910 (jim_thompson5910):

Similar Example The price of a notebook has dropped to 3.30 today. Yesterday's price was 3.75 . Find the percentage decrease. Percent Decrease = [(Old - New)/Old] * 100 Percent Decrease = [(3.75-3.30)/3.75] * 100 Percent Decrease = [0.45/3.75] * 100 Percent Decrease = 0.12 * 100 Percent Decrease = 12% This is NOT the answer to your question as it's just an example

OpenStudy (anonymous):

thank you for your help

jimthompson5910 (jim_thompson5910):

yw